# Usage Governance Advisor: from Intent to AI Governance

### v1

@@ Introduction
The following section evaluates two primary stages of the Usage Governance Pipeline.
@@
The KG is populated by the processing of technical documents about AI models, obtaining related entities (triples) relevant to the ontology previously described. We evaluate different approaches using generative AI against manually obtained ground truth from the granite-8b-code-base-4k AI model card in Hugging Face (IBM 2024). The generated KG is then measured against the ground truth by considering entity/relationship triples and calculating how many of the ground truth triples exist in the prediction (recall) and how many of the predict triples exist in the ground truth (precision). We also provide the F1 metric relating the two for completeness. This F1 metric is computed following the same approach as in (Mihindukulasooriya et al. 2023). Results are provided in Table 1.
@@
In our case, we assume that for two triples to match, the two related entity types must match, the entity labels (names) must also match and the relationship is implicit from the ontology; i.e., it does not matter what the relationship type is, only that the two entities in the triple are related. As an example, consider that an “AI Model” entity is related to a “License” entity; independently of what the process obtains as a relationship type, we assume that the triple refers to the license of said AI model.
@@
Preliminary results showed that having an exact match for triples in the ground truth and triples generated by an LLM is too strict a metric. As an example, consider the following: A triple [’ibm research’, ’organization’, ’granite-8b-code-base-4k’, ’aimodel’] does not match [’granite-8b-code-base-4k’, ’aimodel’, ’ibm research’, ’organization’], although for our purposes they both contain the exact same information. To support these cases, we extend the pipeline with an LLM-as-a-judge solution using granite-3-8b-instruct which improved results as shown in Table 1. Accumulating results over multiple runs for the same document improves the results for the LangChain APIs (LLMGraphTransformer and GraphRag), especially for recall. The process was run up to 15 times, however, performance reached the upper limit within 2-4 passes. Granite-3-8b-instruct as Judge recovered the most information, and was correct around 80% of the time.
Entity/Relationship Extraction Runs P R F1
LLMGraphTransformer 1 0.25 0.36 0.29
with mixtral-8x7b-instruct-v01
3 0.26 0.52 0.35
+ Judge 4 0.70 0.76 0.73
GraphRag with 1 0.27 0.4 0.32
mixtral-8x7b-instruct-v01
2 0.30 0.6 0.40
+ Judge 4 0.71 0.76 0.73
RAG with 1 0.25 0.16 0.19
mixtral-8x7b-instruct-v01
+ Judge 1 0.75 0.6 0.66
Custom prompt with 1 0.17 0.48 0.25
mixtral-8x7b-instruct-v01
+ Judge 1 0.77 0.88 0.82
Custom prompt with 1 0.20 0.64 0.31
granite-3-8b-instruct
+ Judge 1 0.76 0.84 0.80
Overfitted prompt 1 0.92 0.92 0.92
@@
Table 1: Precision (P), Recall (R) and F1 metric for the tuples obtained with different methods.
@@
We evaluated the effectiveness of the Chain-of-Thought (CoT) approach in auto-assisting questionnaire completion with synthetically generated user intents. We generated 42 artificial user intents and corresponding answers using an LLM. These answers were validated by human-annotated ground truth. Accuracy is used to measure binary and dropdown questions. To evaluate free-form questions, the answers generated by CoT were condensed to bullet points to better support accuracy measures and compared with the ground truth.
Figure 6 provides an illustrative example of user intent, response, and bullet-point summary generated by the CoT approach and consequent user summarization. Table 2 shows a subset of the questions used to assess the effectiveness of CoT approach. For each question, we provided few-shot or Chain-of-Thought reasoning examples to facilitate consistent response generation, as shown in Figure 7.
@@
Question type Zero-shot Few-shot/CoT (1 choice) Few-shot/CoT (user choice)
@@
Effective AI governance requires a multifaceted approach that considers diverse stakeholder perspectives (service providers, users, auditors) to mitigate risks. Governance rules are inherently dynamic, subject to frequent updates and new models emerging in the market, making it challenging to ensure compliance. This paper explores the potential of knowledge graphs as a solution for governing AI systems. By integrating new rules, models, and mitigation strategies seamlessly, while maintaining relationships with existing information, KGs can provide a comprehensive framework for managing AI governance complexities. This paper presents an essential first step in addressing regulatory and compliance requirements for deploying AI models into production, supporting alignment with evolving laws and regulations. Real-time monitoring and auditing of AI system performance are also crucial to identify risks associated with AI failures or misuse. To further enhance the framework, future work will focus on developing a robust mechanism for continuously auditing and verifying AI system performance through integration of real-time monitoring capabilities. Additionally, we aim to incorporate policy guidelines from stakeholders directly in the knowledge base, enabling a more tailored approach for AI governance. By addressing these aspects, the framework can provide a comprehensive and proactive approach to AI governance, ensuring responsible deployment and continuous improvement of AI systems.

### v2

@@ Introduction
We evaluate two foundational workflows used in the Usage Governance Advisor that themselves use generative AI.
@@
The KG is populated by extracting entities/relationship in the ontology from technical documentation. The only way to scale this approach is through the use of highly generic fully automated ingestion processes. We evaluate different approaches using generative AI against manually obtained ground truth from the granite-8b-code-base-4k AI model card in Hugging Face (IBM 2024). The generated KG is then measured against the ground truth by considering entity/relationship triples and calculating how many of the ground truth triples exist in the prediction (recall) and how many of the predicted triples exist in the ground truth (precision). Note that precision can be reduced by both incorrect triples, i.e hallucinations, as well as correct but irrelevant ones.
We provide the F1 metric relating the two for completeness. This is the same approach used in (Mihindukulasooriya et al. 2023).
@@
We assume that for two triples to match, the two related entity types must match, the entity labels (names) must also match and the relationship is implicit from the ontology; i.e., it does not matter what the relationship type is, only that the two entities in the triple are related. For example, if an AI Model entity is related to a License entity; then independently of whether the extraction process labeled the relationship is has, released-under, uses etc. we assume that the triple refers to the license of said AI model.
Results are provided in Table 1 for a selection of different extraction methods. One can see that running the process multiple time increases the F1 score due to the random nature of the LLM entity identification, i.e. additional entities are correctly identified in subsequent runs and incrementally added to the graph. The process was run up to 15 times, however, performance reached the upper limit within 2-4 passes, which are reported in the table.
@@
Preliminary results showed that having an exact match for triples in the ground truth and triples generated by an LLM is too strict a metric. As an example, consider the following: A triple [’ibm research’, ’organization’, ’granite-8b-code-base-4k’, ’aimodel’] does not match [’granite-8b-code-base-4k’, ’aimodel’, ’ibm research’, ’organization’], although only the direction of the relationship is distinct. To support these cases, we extend the pipeline with an LLM-as-a-judge solution using granite-3-8b-instruct which improved results as shown in Table 1. The judge improve recall and precision by identifying matches that a simple lexical comparison would miss.
Granite-3-8b-instruct as Judge recovered the most information, and was correct around 80% of the time. Note that other than the ontology itself the extraction process is independent of the nature of the entities and relationships, allowing it be reused even as the ontology evolves. In order to test an upper-bound on the accuracy of the extraction process we also used a set of highly specific prompts, labelled as Overfitted. These do better than all the generic approaches but only by about 8% than the best motivating the usability of the generic approach.
The use of generative AI currently does not allow 100% certainty about the validity of the generated information. In our view there should be a measure of certainty given about extracted facts with suitable links to evidence that a human could check and verify if the certainity is below an acceptable threshold. Measuring the accuracy of model utterances is an active area of AI research.
Entity/Relationship Extraction Match Runs P R F1
LLMGraphTransformer Exact 1 0.25 0.36 0.29
with mixtral-8x7b-instruct-v01 Exact 3 0.26 0.52 0.35
Judge 4 0.70 0.76 0.73
GraphRag with Exact 1 0.27 0.4 0.32
mixtral-8x7b-instruct-v01 Exact 2 0.30 0.6 0.40
Judge 4 0.71 0.76 0.73
RAG with Exact 1 0.25 0.16 0.19
mixtral-8x7b-instruct-v01 Judge 1 0.75 0.6 0.66
Custom prompt with Exact 1 0.17 0.48 0.25
mixtral-8x7b-instruct-v01 Judge 1 0.77 0.88 0.82
Custom prompt with Exact 1 0.20 0.64 0.31
granite-3-8b-instruct Judge 1 0.76 0.84 0.80
Overfitted prompt Exact 1 0.61 0.64 0.62
Judge 1 0.92 0.92 0.92
@@
Table 1: Precision (P), Recall (R) and F1 metric for the triples obtained with different methods.
@@
We evaluated the effectiveness of the Chain-of-Thought (CoT) approach in auto-assisting questionnaire completion with synthetically generated user intents. We generated 42 artificial user intents and corresponding answers using an LLM. These answers were validated by human-annotated ground truth. Accuracy is used to measure binary and dropdown questions. To evaluate free-form questions, the answers generated by CoT were condensed to bullet points to better support accuracy measures and compared with the ground truth. Figure 6 provides an illustrative example of user intent, response, and bullet-point summary generated by the CoT approach and consequent user summarization. Table 2 shows a subset of the questions used to assess the effectiveness of CoT approach. For each question, we provided few-shot or Chain-of-Thought reasoning examples to facilitate consistent response generation (see Figure 7).
@@
Question type Zero-shot Few-shot/CoT Few-shot/CoT
(1 choice) (user choice)
@@
Effective AI governance requires a multifaceted approach that considers diverse stakeholder perspectives (service providers, users, auditors) to mitigate risks. Governance rules are inherently dynamic and subject to frequent updates. In addition, new models are continuously emerging in the market. This makes it challenging to ensure that all risks have been appropriately considered. This paper explores the potential of knowledge graphs as a solution for governing AI systems. By integrating new rules, models, and mitigation strategies seamlessly, while maintaining relationships with existing information, KGs can provide a comprehensive framework for managing AI governance complexities. We present an essential first step in addressing regulatory and compliance requirements for deploying AI models into production, supporting alignment with evolving laws and regulations. Real-time monitoring and auditing of AI system performance are also crucial to identify risks associated with AI failures or misuse. To further enhance the framework, future work will focus on developing a robust mechanism for continuously auditing and verifying AI system performance through integration of real-time monitoring capabilities. Additionally, we aim to incorporate policy guidelines from stakeholders directly into the knowledge base, enabling a more tailored approach to AI governance. By addressing these aspects, the framework can provide a comprehensive and proactive approach to AI governance, ensuring responsible deployment and continuous improvement of AI systems.